### v1

# Dynamical scheme for computing the mass parameter of a system in a medium

Abstract
We present a new method for extracting a mass parameter using time-dependent density functional theory for an arbitrary physical system, provided the adiabatic limit is achievable. This approach works for collective variables also in the presence of a medium, in particular for the nuclei interacting with a neutron background. We apply the method to extract mass parameters of impurities in the neutron star crust, like their inertial masses and quadrupole mass parameters. The extracted mass parameters at various depths of the inner crust are compared with other methods, including the hydrodynamic approach. The presented method opens avenues for the construction of an effective model of neutron star crust grounded in microscopic calculations.
I Introduction
Effective models are indispensable tools in physics that inevitably emerge for complex systems. With their help, one can connect a physical phenomenon taking place across various scales. Whenever we change degrees of freedom and construct a new effective model operating at larger length scales, input from the underlying model is required. In this way, we obtain a hierarchy of models spanning from macroscopic approaches through mesoscopic ones up to microscopic ones. As an example, relevant to this paper, let us consider the problem of constructing an effective model of a rotating neutron star. At the largest scales, it can be modeled by means of hydrodynamical approaches Graber et al. (2017); Shibata et al. (2017). To explain observed phenomena of neutron star glitches (sudden decreases of rotation period), one needs to revert to a multifluid picture, with the extra assumption that one of the fluids is in a superfluid state Haskell and Melatos (2015); Khomenko and Haskell (2018). Besides the equation of state, such models require various macroscopic parameters to be provided, like viscosities, conductivities, and mutual friction parameters between the fluid components. Going down in the hierarchy, by increasing the resolution, we enter the regime of mesoscopic models, where we start to resolve quantized vortices present in the superfluid component, which coexists with the so-called impurities created by the other fluid component Howitt et al. (2020); Cheunchitra et al. (2024). Precisely, neutron matter forms the superfluid component, filled out with vortices, while protons form various structures, which for the majority of the inner crust have spherical shapes. Electrons, which are essential for charge neutrality, are well approximated by their uniform distribution. At this level, one usually refers to the Vortex Filament Model (VFM), which was originally developed for modeling vortices in superfluid helium Tsubota et al. (2013). It is a quasiclassical model, where one writes Newton’s equations for quantum vortices (presented as filaments in 3D or points in 2D) that interact with all other elements of the system. The drag force exerted on vortices by the impurities is the origin of the mutual force that appears in the hydrodynamical model. Increasing further the resolution, we start to resolve the building blocks of the neutron’s star matter: neutrons, protons, and electrons (and possibly other elementary particles deep in the star’s core). At this level, one needs to revert to the microscopic description grounded in many-body quantum mechanics. The quantized vortices turn out to be a collective flow of neutrons around some line in space, while impurities correspond to regions where protons (together with neutrons) concentrate into various structures Negele and Vautherin (1973); J. M. Pearson, N. Chamel, A. Y. Potekhin, A. F. Fantina, C. Ducoin, A. K. Dutta and S. Goriely (2018), usually into nuclei-like objects, but more complex structures are also allowed Caplan and Horowitz (2017). The effective interaction between the flow of neutrons and the bound state of protons is the origin of the drag force that enters the mesoscopic models.
The process of extracting effective model parameters from the underlying approach is a complex problem. Here, we focus on the method of extracting one of the most fundamental quantities that enters into effective models, namely mass parameters. Basically, each effective model contains a term that characterizes the inertia of the selected degree of freedom $Q$. Schematically, the effective Hamiltonian can be written
$H_{Q}\approx\frac{M_{Q}\dot{Q}^{2}}{2}+V(Q),$ (1)
where $Q$ is the selected collective variable. Such a form is achievable once the adiabatic limit can be recovered in the system, which means that the terms describing couplings to other degrees of freedom can be neglected Baranger and Vénéroni (1978). The first term is the kinetic term that is quantified by the mass parameter $M_{Q}$, and $V$ stands for the potential energy. The choice of the collective variable is motivated by physical intuition, although a more precise definition is also possible. In the latter case, one searches for the collective path (or space), which is maximally decoupled from the other degrees of freedom (see Washiyama et al. (2024); Wen and Nakatsukasa (2022); Nakatsukasa et al. (2016) and references therein). This can be important in the case where the choice of collective variables is not obvious, as in the case of exotically deformed nuclei forming the pasta phase. Whatever method is applied to identify $Q$, it implicitly assumes that the adiabatic limit can be achieved, in which the motion of the system is governed by Hamiltonian (1).
In this paper, we present a general-purpose method for extracting mass parameters for nuclear systems by means of the time-dependent density functional theory approach (TDDFT). While there are methods for extracting the effective mass parameters, most of them rely on various assumptions, and they are typically designed to be applicable for extracting the mass parameters in a vacuum (see, e.g., the review paper Li and Vretenar (2022)). The presented approach is free of these limitations, and we will demonstrate it for the nuclear impurity immersed in superfluid neutron liquid, i.e., for the conditions relevant to the neutron star crust. The calculations will be delivered for mass parameters associated with translational motion and quadrupole vibrations. The selected collective variables are schematically presented in Fig. 1.
Figure 1: Scenario considered in this work: a nucleus 40Zr immersed in a superfluid neutron matter. The simplest types of collective modes that nuclei can execute are: (a) center of mass and (b) quadrupole moment oscillations. From the time evolution of the collective degrees of freedom associated with the mode, we can fit the parameters of interest, such as the mass parameter and dissipation coefficients.
The article is structured as follows: in Sec. II, we introduce the aim and context of this work. In Sec. III, we apply the method to two collective variables: the center of mass and the quadrupole moment. Finally, in Sec. IV, we conclude the paper.
II Method
To make the problem more concrete, let us consider a Wigner crystal of neutron-rich nuclei in the crust of a neutron star. These nuclei can oscillate within the crystal, as illustrated in Fig. 1(a). The collective variable $Q$ can then be associated with the impurity’s position $\bm{R}_{\textrm{CM}}$, which, in a microscopic description, corresponds to the center of mass of the protons. The potential energy in this system is primarily due to Coulomb interactions between impurities. Even in this relatively simple setting, determining the effective mass of an impurity immersed in neutron matter remains a challenging task. A variety of approaches have been developed to address this problem Magierski (2004); Magierski and Bulgac (2004); Martin and Urban (2016); N. Chamel, D. Page and S. Reddy (2013); Pęcak et al. (2024). The Wigner crystal is also expected to be threaded by quantum vortices. Work G. Wlazłowski, K. Sekizawa, P. Magierski, A. Bulgac and M. McNeil Forbes (2016) demonstrated that the interaction between an impurity and a vortex can induce a quadrupole deformation of the impurity, Fig. 1(b). This suggests that the quadrupole moment $Q_{20}$ could also serve as a relevant collective variable in the effective model. It is also possible that higher-order nuclear multipole moments play a significant role.
The shape of the impurity can be parametrized using an expansion in spherical harmonics $Y_{\lambda\mu}(\theta,\phi)$
$R(\theta,\phi)=R_{0}\left(1+\sum_{\lambda=0}^{\infty}\sum_{\mu=-\lambda}^{\lambda}\alpha_{\lambda\mu}Y_{\lambda\mu}(\theta,\phi)\right).$ (2)
where $R_{0}$ is the radius of the undeformed impurity, and each deformation parameter $\alpha_{\lambda\mu}$ is associated with a corresponding mass parameter. In this expansion, $\lambda=1$ corresponds to the dipole mode, which relates to the impurity’s position, while $\lambda=2$ corresponds to the quadrupole mode. These two modes have been identified as relevant collective variables for the mesoscopic model and are therefore the focus of investigation in this work.
In order to investigate the dynamics of the many-body system, we will consequently use an approach based on the density functional theory (DFT). In the case of nuclear systems, it offers a computationally tractable method, where the energy functional is constructed using various types of densities relevant to the problem. In recent decades, it has become one of the standard theoretical tools to study medium and heavy nuclei Nakatsukasa et al. (2016); Colò (2020); Yang and Piekarewicz (2020). The heart of each DFT method is an energy functional
$E(t)=E[\rho(\bm{r},t),\ldots],$ (3)
which maps the density of particles $\rho$ at a given position $\bm{r}$ and time $t$ into instantaneous energy (generally, the functional can also depend on other densities marked by dots). We assume that the density is also sufficient to define a collective variable
$Q(t)=Q[\rho(\bm{r},t)].$ (4)
The aim is to construct the effective Hamiltonian for the collective variable, which means that we would like to approximate the energy (3) by a form given by Eq. (1). Suppose we have at hand a microscopic method that can provide time evolution of the density $\rho(\bm{r},t)$, and related quantities, like $Q(t)$. If we restrict our considerations to cases where the collective variable exhibits small fluctuations $\delta Q(t)=Q(t)-Q_{0}$ with respect to its equilibrium value $Q_{0}$, then the effective potential can be well approximated as
$V(Q)\approx\frac{1}{2}M_{Q}\omega_{Q}^{2}\left(Q(t)-Q_{0}\right)^{2}.$ (5)
In such a case, the recorded dynamics of $Q(t)$ should follow harmonic oscillations with frequency $\omega_{Q}$. The measured frequency provides us with information about the mode’s energy $E_{Q}=\hbar\omega_{Q}$. However, it is insufficient to deliver information about the mass parameter $M_{Q}$. But we can consider dynamics arising from a modified functional, supplemented by an additional harmonic potential that couples to the collective variable
$E^{\prime}(t)=E(t)+\frac{1}{2}k\left(Q(t)-Q_{0}\right)^{2}.$ (6)
Upon this modification, we can rearrange the terms so that the effective potential takes the form
$V^{\prime}(Q)\approx\frac{1}{2}M_{Q}\omega^{2}(k)\left(Q(t)-Q_{0}\right)^{2},$ (7)
where
$\omega^{2}(k)=\omega_{Q}^{2}+\frac{k}{M_{Q}},$ (8)
and the collective variable will oscillate with modified frequency $\omega(k)$ that is sensitive to the mass parameter $M_{Q}$. The procedure for extracting the mass parameter is then as follows:
  1. 1.
    Solve time-dependent equations of motion arising from the modified energy functional $E^{\prime}$, Eq. (6), for different values of parameter $k$ and for the case corresponding to the small amplitude oscillations in the variable $Q(t)$.
  2. 2.
    Extract frequency of oscillations $\omega(k)$ and fit to it formula (8) with respect of parameters $\omega_{Q}$ and $M_{Q}$.
The procedure provides the mode’s energy $\omega_{Q}$ and the associated mass parameter $M_{Q}$.
It is worth mentioning that there are other methods to extract collective mass parameters that are typically defined in a vacuum Li and Vretenar (2022). As a result, no existing method works for nuclei in a medium, such as the inner crust of a neutron star. One example is the adiabatic approximation to the time-dependent Hartree-Fock-Bogoliubov approach (ATDHFB), commonly used to analyze large amplitude collective motion Baran et al. (2011). The fundamental idea behind this method is that the collective motion of the whole system is slow compared to the independent movement of each of the nucleons that the system consists of. This approach has various modifications, including the popular cranking approximation Girod and Grammaticos (1979). All these approaches extract the mass parameter from quasiparticle wave function, corresponding to static solutions of time-independent HFB equations ($i\hbar\partial/\partial t\rightarrow E_{\eta,q}$ in Eq. (9)). Such methods encounter difficulties in cases of more complex problems, like nuclei immersed in the neutron matter. In such a case, one cannot easily split the quasiparticle wave functions into two sets, where one represents nucleons belonging to nuclei and the other represents nucleons belonging to the background. The approach presented here is free of these difficulties, as it relies only on the observables constructed from densities. Therefore, the method is universal and can be applied to the region of the inner crust of a neutron star to extract the effective mass parameters associated with various collective degrees of motion.
Equations of motion that arise from modern energy functionals for nuclear systems typically have a structure formally equivalent to time-dependent Hartree-Fock-Bogoliubov (HFB) equations (position and time dependence have been omitted for brevity)
$i\hbar\frac{\partial}{\partial t}\begin{pmatrix}u_{\eta,q\uparrow}\\ u_{\eta,q\downarrow}\\ v_{\eta,q\uparrow}\\ v_{\eta,q\downarrow}\end{pmatrix}=\begin{pmatrix}h_{q\uparrow\uparrow}&h_{q\uparrow\downarrow}&0&\Delta_{q}\\ h_{q\downarrow\uparrow}&h_{q\downarrow\downarrow}&-\Delta_{q}&0\\ 0&-\Delta_{q}^{*}&-h^{*}_{q\uparrow\uparrow}&-h^{*}_{q\uparrow\downarrow}\\ \Delta_{q}^{*}&0&-h^{*}_{q\downarrow\uparrow}&-h^{*}_{q\downarrow\downarrow}\end{pmatrix}\begin{pmatrix}u_{\eta,q\uparrow}\\ u_{\eta,q\downarrow}\\ v_{\eta,q\uparrow}\\ v_{\eta,q\downarrow}\end{pmatrix}.$ (9)
Index $q\in\{n,p\}$ stands for isopin labeling either neutron or proton. Arrows indicate the spin, and $\eta$ is the quantum number labeling the wave functions. The diagonal part of the Hamiltonian matrix contains mean-field contribution, defined via functional derivative $U_{q}(\bm{r},t)={\delta E}/{\delta\rho_{q}}$. Thus, the equations of motion arising from the modified functional (6) account for simple modification ($\sigma\in\{\uparrow,\downarrow\}$)
$h_{q\sigma\sigma}\rightarrow h_{q\sigma\sigma}+k\left(Q(t)-Q_{0}\right)\frac{\delta Q}{\delta\rho_{q}}.$ (10)
The method described above is generic and does not depend on specific functional details. In this study, we use the Brussels-Montreal class of Skyrme-type functionals, which are particularly well suited for studies in the context of neutron stars. These functionals are developed using not only experimental data biased on finite-size nuclei but also many-body results Cao et al. (2006) that account for infinite nuclear matter. In this way, the bulk properties are incorporated into the model. The Brussels-Montreal-Skyrme functional has a generic form:
[…]
with
$\mathcal{E}_{\textrm{BSk}}[\rho_{q},\nabla\rho_{q},\tau_{q},\bm{j}_{q},\nu_{q}]=$
$\sum_{q}\frac{\hbar^{2}}{2m_{q}}\tau_{q}+\mathcal{E}_{\rho}+\mathcal{E}_{\Delta\rho}+\mathcal{E}_{\tau}+\mathcal{E}_{\pi}+\mathcal{E}_{C}.$ (12)
In this functional, one can distinguish the kinetic term ($m_{q}$ being the bare mass of a given nucleon), $\mathcal{E}_{\rho}$ describes the properties of uniform nuclear matter at certain density, $\mathcal{E}_{\Delta\rho}$ accounts for corrections coming from nonuniformities (so-called gradient terms), $\mathcal{E}_{\tau}$ contributes to the effective mass of nucleon, $\mathcal{E}_{\pi}$ is connected with pairing correlations and $\mathcal{E}_{C}$ refers to Coulomb interaction. The explicit form of these terms is given in Ref. Pęcak et al. (2024). In the context of this work, it is sufficient to mention that they depend on densities, which are defined through Bogoliubov amplitudes
$\rho_{q}(\bm{r},t)$ $=\sum_{\eta\sigma}\left|v_{\eta,q\sigma}(\bm{r},t)\right|^{2}$ (13)
$\tau_{q}(\bm{r},t)$ $=\sum_{\eta\sigma}\left|\bm{\nabla}v_{\eta,q\sigma}(\bm{r},t)\right|^{2},$ (14)
$\nu_{q}(\bm{r},t)$ $=2\sum_{\eta}\left[u_{\eta,q\uparrow}(\bm{r},t)v^{*}_{\eta,q\downarrow}(\bm{r},t)-u_{\eta,q\downarrow}(\bm{r},t)v^{*}_{\eta,q\uparrow}(\bm{r},t)\right],$ (15)
$\bm{j}_{q}(\bm{r},t)$ $=\sum_{\eta\sigma}\mathrm{Im}\left[v_{\eta,q\sigma}(\bm{r},t)\bm{\nabla}v^{*}_{\eta,q\sigma}(\bm{r},t)\right].$ (16)
The densities in the above equations are the particle number density ($\rho_{q}$), kinetic density ($\tau_{q}$), anomalous density ($\nu_{q}$), and current density ($\bm{j}_{q}$). The summation is taken over the quasiparticle states whose energy $E_{\eta,q}$ is smaller than the cutoff energy $0<E_{\eta,q}<E_{\textrm{cut}}$. All elements in the HFB matrix are expressed via suitable functional derivatives of the energy density functional with respect to these densities. For example, the diagonal terms are
$h_{q\sigma\sigma}=-\bm{\nabla}\frac{\delta E}{\delta\tau_{q}}\bm{\nabla}+\frac{\delta E}{\delta\rho_{q}}-\frac{i}{2}\left\{\frac{\delta E}{\delta\bm{j}_{q}},\bm{\nabla}\right\},\$ (17)
[…]
In this work, we will be considering oscillations of a collective variable only around its ground state value. It means that $E(t)$ from Eq. (3) is equivalent to $E_{\textrm{BSk}}(t)$. However, in general, one can consider oscillations around arbitrary values of the collective variable by considering a constrained functional, like
$E=E_{\textrm{BSk}}-\lambda_{Q}\left(Q-Q_{0}\right),$ (18)
where $\lambda_{Q}$ is the Lagrange multiplier assuring that the functional has a minimum at the requested point in the collective space. Next, the oscillations may be damped, which reflects the finite lifetime of various modes. To improve the quality of the mass parameter determination, we use the model of the damped harmonic oscillator when extracting the frequency of the oscillations:
$Q(t)=Ae^{-\gamma t}\sin(\omega_{\textrm{d}}t+\varphi),$ (19)
where $\gamma$ is the damping coefficient, and $A$ the initial amplitude of the perturbation. Note that we take into account the correction to the frequency of damped oscillations $\omega_{\textrm{d}}$
$\omega^{2}=\omega_{\textrm{d}}^{2}+\gamma^{2}.$ (20)
The undamped $\omega$ frequency is next used in Eq. (8) for obtaining the mass parameter $M_{Q}$.
III Results
We apply the method to a specific nuclear system: the inner crust of a neutron star. In this region, the crystalline lattice of bound nuclei is permeated by superfluid neutrons. This configuration naturally leads to a deeper mean-field proton potential and a higher concentration of neutrons near protons. Since these exotic neutron-rich nuclei lack a well-defined radius and have diffuse surfaces, defining neutron-based observables presents methodological challenges. However, we assume that the overall behavior of the nucleus can be characterized by tracing the proton component, which serves as the fundamental basis for measuring collective observables.
III.1 Center of mass oscillations
The inner crust of a neutron star is an example of a many-body polaron system Baroni et al. (2024), where a bound system (nucleus) is immersed in and interacts with a bath of superfluid neutrons. As extensively discussed in Pęcak et al. (2024), extracting the effective mass of such a self-bound object is a very complex problem. On the other hand, understanding the effective mass of nuclei in the inner crust is crucial for estimating, for example, the crystalline lattice phonon spectrum or thermal and electric conductivities Baiko et al. (2001); Gnedin et al. (2001).
The effective mass of the impurity describes its ability to move in the medium. In principle, it can be spatially dependent, and in an anisotropic medium, it can form a tensor. Since our setup is spatially uniform, and we neglect the potential effect of the bands from the crystalline lattice Chamel (2005), the effective mass is a scalar. We will consider oscillations of the center of mass along $z$ direction
[…]
$\begin{split}U_{\textrm{p}}=\frac{\delta E}{\delta\rho_{p}}=U_{\textrm{p},\textrm{BSk}}+k\left(Z_{\textrm{CM}}(t)-Z_{0}\right)\frac{z}{N_{p}}.\end{split}$ (22)
One can recognize that the new term in the mean field corresponds to a force (computed as the gradient of the potential) which is constant in space, but the strength is time-dependent. It becomes evident that such a force can move a cluster of protons in space, but cannot deform their density distribution.
In order to examine the system described above, we did systematic studies for average nuclear densities $\bar{\rho}$ ranging from 0.002 $0.002\textrm{~{}fm}^{-3}$ to 0.051 $0.051\textrm{~{}fm}^{-3}$, where the average density is defined as $\bar{\rho}=(N_{n}+N_{p})/V_{WS}$, ie., number of nucleons (protons and neutrons) per Wigner-Seitz cell volume J. M. Pearson, N. Chamel, A. Y. Potekhin, A. F. Fantina, C. Ducoin, A. K. Dutta and S. Goriely (2018). We have generated initial states, as shown in Fig. 1, by means of solving the static variant of problem Eq. (9). The problem was solved on a spatial mesh of size $N_{x}\times N_{y}\times N_{z}=32\times 32\times 64$ with lattice spacing $\Delta x=\Delta y=\Delta z=1.25$ fm. The associated cutoff energy is fixed through the value of resolution to $E_{\mathrm{cut}}=\hbar^{2}\pi^{2}/(2m_{n}\Delta x^{2})\approx 130$ MeV. In the computation, we have used W-BSK Toolkit, described extensively in Ref. Pęcak et al. (2024), and available as an open-source package via webpage WBS . As a side remark, we note that the toolkit neglects contributions arising from the spin-orbit part of the nuclear interaction, which in turn allows the execution of simulations in sufficiently large volumes relevant to nuclear problems in the context of neutron star studies. Next, we have performed time-dependent simulations with modified mean-field potential Eq. (22) with the initial displacement $Z_{0}=1$fm.
Figure 2: a) Position of the center of mass of protons as a function of time for different values of the $k$ parameter. The considered system has average nuclear density $\bar{\rho}=0.015\textrm{~{}fm}^{-3}$. b) The extracted frequency of oscillations as a function of the $k$ parameter. The dashed line is the fit of Eq. (8).
We applied the fitting procedure to extract the effective parameters of the impurity. Fig. 2 shows the center of mass oscillations for different values of control parameter $k$, for selected nuclear density $\bar{\rho}=0.015\textrm{~{}fm}^{-3}$. As expected, the extracted frequencies of oscillations follow Eq. (8). One can clearly see that the intercept in this case is $\omega_{Z_{\textrm{CM}}}=0$. This feature of the center of mass collective degree of freedom is also expected. Since the Hamiltonian is spatially symmetric, i.e., it is invariant under the operation $\bm{r}\rightarrow\bm{r}+\bm{r}_{0}$, shifting the nucleus does not change the energy of the system. Nevertheless, the translational symmetry is broken: the nuclear cluster can form at any place, which is set randomly. The $\omega_{Z_{\textrm{CM}}}=0$ mode is then analogous to a zero-energy Goldstone mode associated with the breaking of translational symmetry. This mode does not represent oscillation but rather a shift with a constant velocity that moves the ground state across all degenerate quantum states that are also solutions of the Hamiltonian.
Figure 3: The effective mass $M_{\textrm{CM}}$ (in units of neutron mass $m_{n}$) as a function of mean nuclear density $\bar{\rho}$ extracted for small amplitude $A=1$ fm (orangle triangles) and large amplitude $A=1$ fm (green squares) oscillations. We compare the results of the proposed method (orange triangles) with the effective mass obtained recently using a different dynamical scheme (blue diamonds) Pęcak et al. (2024).
The extraction of the effective mass of the nucleus as a function of the inner crust mean nuclear density $\bar{\rho}$ is summarized in Fig. 3. The results are benchmarked against independent calculations performed within the same numerical framework Pęcak et al. (2024) but based on a different scheme. This benchmark scheme is based on the measurement of acceleration once a constant external force is applied to the nucleus. It is specific to the center-of-mass collective variable and cannot be extended to other collective variables, a limitation that our current approach overcomes. The results are consistent within numerical uncertainties, which remain below 5%.
Figure 4: a) The damped oscillations of the center of mass of protons for selected density $\bar{\rho}=0.019\textrm{~{}fm}^{-3}$ (we show that point in panel c), and b) associated absolute value of velocity as a function of time. The velocity is compared to Landau’s velocity. In panel c), we show the extracted damping coefficient $\gamma_{\textrm{CM}}$ as a function of average nuclear density.
The extraction of the mass parameter is valid in the limit of small-amplitude oscillations. In this regime, no damping is observed (see Fig. 2a), as the nucleus moves through the superfluid without resistance. This behavior reflects a fundamental property of superfluids: as long as the velocity remains below a critical threshold $v_{\textrm{L}}$, the system does not exert a drag force on the moving object. To explore dissipative effects, the numerical experiment can be readily adjusted to achieve higher maximum velocities by increasing the oscillation amplitude $A$. The resulting maximum velocity is proportional to $\max\left[V_{\textrm{CM}}\right]=A\omega_{d}$, where $V_{\textrm{CM}}(t)=\dot{Z}_{\textrm{CM}}$. We have repeated the simulations with an increased value of amplitude $A=10$ fm, which is of the same order as the radius of the impurity. For the new setup, the maximum velocity exceeds this so-called Landau velocity
$v_{\textrm{L}}=\frac{\Delta_{n}}{\hbar{k_{\mathrm{F}}}_{n}},$ (23)
where $\Delta_{n}$ and ${k_{\mathrm{F}}}_{n}$ are the bulk values of the pairing field and the Fermi wave-vector of neutrons. Note that this quantity is defined for a uniform system, and it is not obvious how well it works for such a non-uniform system as we study. For increased amplitude, we observed damped oscillations as shown in Fig. 4a). From fitting to the Eq. (19) we can accurately reproduce the behaviour of damped oscillations (orange solid line) and extract the damping parameter $\gamma_{\textrm{CM}}$. We investigate the absolute value of the velocity of the impurity in Fig. 4b). One can clearly connect the damping with exceeding the Landau velocity by the impurity. The damped oscillator model starts to underestimate the oscillations’ amplitude once the maximum velocity drops significantly below the Landau velocity. A similar mechanism was observed in recent work Pęcak et al. (2024) where the breaking of Cooper pairs was one of the mechanisms of energy dissipation. The last panel of Fig. 4c) shows the damping parameter $\gamma_{\textrm{CM}}$ as a function of average nuclear density in the inner crust. This is a monotonically increasing function of the density. In the limit of low density $\bar{\rho}\rightarrow 0$, it corresponds to a nucleus oscillating in the vacuum. Obviously, in this regime, there are no degrees of freedom where the energy can be dissipated, thus no damping is present. Once the medium gets denser, we see that the damping coefficient increases as well. This effect comes from the investigation of Eq. (23) where the pairing field $\Delta_{n}$ vanishes in high densities Cao et al. (2006), while the wave vector in the denominator $k_{\mathrm{F}}$ grows. Note that in this work, we consider only the ${}^{1}S_{0}$ pairing channel, and neglect other channels for the pairing, such as higher angular momentum channels ${}^{3}P_{1}$, which is believed to dominate in higher densities.
The case of oscillations with amplitude $A=10$ fm cannot be considered small, as the amplitude becomes comparable to, or even exceeds, the size of the protonic cluster itself, whose radius ranges from $5$ to $10$ fm depending on the density. Nevertheless, extracting the mass parameter from these data still yields reasonable results, see square markers in Fig. 3. Although the method consistently overestimates $M_{\textrm{CM}}$ by approximately $10-15$%, the qualitative behavior is well captured. This discrepancy can be viewed as an estimate of the upper bound on the method’s uncertainty.
III.2 Quadrupole moment oscillations
The center-of-mass collective variable of a spherical nucleus provides the simplest example, serving as an ideal benchmark against existing results in the literature. A slightly more complicated case involves the collective variable associated with nuclear deformation. As the next example, we therefore consider axial quadrupole oscillations of nuclei immersed in a sea of superfluid neutrons. Collective modes, including quadrupole modes, are gapless, meaning they allow for excitations that can contribute, for example, to the heat capacity and transport properties of the inner crust, features that are vital for modeling the thermal evolution of neutron stars Di Gallo et al. (2011); Urban and Oertel (2015). Hydrodynamic models perform coarse-graining of physical quantities and are valid in the limit of small quasimomenta, making our microscopic approach complementary to such frameworks. Apart from the hydrodynamic approach, quadrupole cluster vibrations in the inner crust of neutron stars have also been studied using the quasiparticle random phase approximation (QRPA), with the main emphasis placed on extracting the mode’s energy Khan et al. (2005); Inakura and Matsuo (2019).
We define the collective variable as follows
[…]
It vanishes for the spherical shape, is negative for oblate, and positive for prolate shapes. Division into nucleus and superfluid neutrons would be artificial. Therefore, to avoid any issues with classifying those neutrons into one belonging to the superfluid medium and those in the cluster, we define again the collective variable using only the proton density. We consider quadrupole oscillations around the spherical shape, initializing the dynamics with a deformed nucleus. Specifically, we prepare initial states with a small nonzero initial value of $Q_{20}(t=0)\approx 100\,\textrm{~{}fm}^{2}$, while the typical deformation of a heavy nucleus is of order $10^{3}-10^{4}\,\textrm{~{}fm}^{2}$. Technically, we achieve this by solving the static problem with a Lagrange constraint as in Eq. (18), within a box of size $N_{x}\times N_{y}\times N_{z}=32\times 32\times 32$ and the lattice spacing $\Delta x=\Delta y=\Delta z=1.25$ fm. Once the initial state is obtained, we perform the time-dependent simulations with a properly adjusted single-particle mean-field potential for protons
$\begin{split}U_{\textrm{p}}=U_{\textrm{p},\textrm{BSk}}+kQ_{20}(t)(3z^{2}-r^{2}).\end{split}$ (25)
Applying such potential results in oscillations between oblate and prolate shapes consecutively.
Figure 5: a) Oscillations of the quadrupole moment $Q_{20}$ as a function of time for different values of the $k$ parameter. The considered system has average nuclear density $\bar{\rho}=0.015\textrm{~{}fm}^{-3}$. b) The extracted frequency of oscillations as a function of the $k$ parameter. The dashed line is the fit of Eq. (8).
Representative results for the selected density $\bar{\rho}=0.015\,\textrm{~{}fm}^{-3}$ are shown in Fig. 5 for various values of the control parameter $k$. In all cases, we observe that the quadrupole mode is damped, even in the limit of small-amplitude oscillations, indicating its finite lifetime. Nevertheless, the time evolution of $Q_{20}(t)$ is well described by a damped harmonic oscillator model, allowing the extraction of the oscillation frequency. The dependence of the frequency on the parameter $k$ follows the expected behavior given by Eq (8). It is worth noting that, unlike the center-of-mass oscillations, the quadrupole mode is a finite-energy excitation, with $\hbar\omega_{Q_{20}}>0$.
Figure 6: a) The energy of the quadrupole mode excitations as a function of the inner crust density $\bar{\rho}$. b) The mass parameter $M_{Q_{20}}$ calculated with TDDFT (green diamonds) and the hydrodynamic approach (red circles).
We extract the oscillation frequencies systematically for various depths of the inner crust. In Fig. 6 we present results for the mode’s energy $\hbar\omega_{Q_{20}}$ and the associated mass parameter $M_{Q_{20}}$ as a function of the mean density $\bar{\rho}$ of the system. The $Q_{20}$ excitation energy decreases with the density. Furthermore, we see that the mass parameter $M_{Q_{20}}$ is also decreasing. This quantity shows how easy it is to deform a nucleus. We observe that the presence of background neutrons diffuses the border of the impurity and allows for deformation at a lower energy cost. It is understandable, as it is known that by increasing density, the protons cease forming spherical clusters in neutron matter but instead evolve into exotic shapes known under the common name nuclear pasta Caplan and Horowitz (2017). The spherical clusters predominantly transform into rod-like structures, which inevitably requires increasing the quadrupole moment to large values at a low energy cost. It is also interesting to note that the simple linear extrapolation of the mass parameter and the mode’s energy points at the density $\approx 0.09\textrm{~{}fm}^{-3}$ as the value around which it vanishes. This density is close to the expected density at which the transition between the crust and the core happens Haensel et al. (2007).
The mass parameter can be estimated with the help of the hydrodynamic approach (see Appendix A for derivation of the formula):
$M_{Q_{20}}=\frac{16\pi m_{n}^{3}}{5M_{\textrm{CM}}^{2}}\rho_{\textrm{out}}(\kappa-1)^{2}\frac{R_{0}}{2\kappa+3},$ (26)
where $\kappa=\rho_{\textrm{in}}/\rho_{\textrm{out}}$, with $\rho_{\textrm{out}}$ being density of neutrons far from the impurity and $\rho_{\textrm{in}}$ being density of protons and neutrons inside the impurity. We compare this prediction with our results in Fig. 6b). We find that for higher densities, the results given by the hydrodynamic method stand in good agreement with the results that we obtained. On the other hand, for densities below 0.03 $0.03\textrm{~{}fm}^{-3}$, the values obtained from hydrodynamics overestimate our predictions, in the low-density limit by a factor of about two. The discrepancies arise, e.g., from the estimation of the radius of impurity — in the hydrodynamical model, we assume a sharp radius of the impurity.
IV Conclusions
We have introduced a new method based on time-dependent density functional theory for extracting mass parameters. This approach is broadly applicable to a wide range of physical systems and collective variables, provided that the variable of interest can be expressed in terms of the density. In principle, this poses no limitation, as the foundational theorems of density functional theory guarantee that all observables can be written as functionals of the density. In practice, however, the explicit functional dependence may not be known for some observables.
The method is well-suited to complex scenarios, such as those encountered in studies of neutron star interiors, where crystalline nuclear structures coexist with superfluid neutrons. Standard techniques often become impractical in such contexts, but the proposed approach offers a promising alternative. As a demonstration, we applied the new method to two cases: extracting the effective mass of nuclear clusters embedded in superfluid neutron matter, and determining the mass parameter associated with quadrupole oscillations. For the former, our results are consistent with earlier findings Pęcak et al. (2024). For the latter, the extracted mass parameter provides new insights into the system’s behavior, characterizing the inertia of an impurity with respect to quadrupole deformations. Notably, the results reveal the increasing significance of this mode in lattice vibrations as the density increases.
The development of new methods grounded in density functional theory lays the foundation for constructing effective models of complex systems such as neutron stars. In particular, combining static and time-dependent approaches appears to be especially powerful. Some quantities, such as the effective potential $V(Q)$ in Eq.(1), are most efficiently extracted using static methods. Others, typically associated with dynamical behavior, are better addressed using time-dependent approaches. Here, we demonstrated the latter by extracting the mass parameter $M_{Q_{20}}$. Together, these complementary methods can provide robust input for collective models of neutron stars Kobyakov and Pethick (2013); Martin and Urban (2014); Di Gallo et al. (2011), which are essential for linking microscopic properties to macroscopic observational data.
A realistic effective model of a rotating neutron star will also need to incorporate quantum vortices. These topological defects are often assumed to be massless Barenghi et al. (2006), although several studies suggest that vortices in Fermi superfluids may in fact carry mass Kopnin and Salomaa (1991); Kopnin and Vinokur (1998); Šimánek (1995); Volovik (1998); Kopnin (2002); Simula (2018). The TDDFT-based method presented here is not suited for determining the effective mass of such defects. However, a recent study has proposed an approach capable of quantifying the mass of quantum vortices using TDDFT Richaud et al. (2024).
Acknowledgements
This work was financially supported by the (Polish) National Science Center Grants No. 2021/40/C/ST2/00072 (DP), 2021/43/B/ST2/01191 (PM, GW). We acknowledge Polish high-performance computing infrastructure PLGrid for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through PLL/2023/04/016476.
Author contributions:
Research planning and construction of the method GW, AZ; numerical calculations: AZ; results analysis: AZ, DP, GW; hydrodynamic model PM, AZ. All authors contributed to the interpretation of the results and manuscript writing.
Data availability:
Data underlying the study and detailed instructions on how to reproduce the results are available via Zenodo repository Zdanowicz et al. (2025).
Appendix A
In this section, the derivation of the formula for the mass parameter within the hydrodynamic approach is presented.
We regard the nuclear matter as an incompressible and irrotational fluid. It implies that there exists a velocity potential $\Phi$ such that $\bm{v}=\nabla\Phi$. We consider impurity of density $\rho_{\textrm{in}}$ and varying with angles radius $R(\theta,\phi)$, in medium of density $\rho_{\textrm{out}}$, see Fig. 7.
Figure 7: Schematic view of a system analyzed within the hydrodynamic approach: impurity with radius $R$ and intrinsic density $\rho_{\textrm{in}}$ moving in medium of density $\rho_{\textrm{out}}$.
We define fields inside ($\Phi_{\textrm{in}}$) and outside ($\Phi_{\textrm{out}}$) of the impurity, and by utilizing spherical symmetry of the considered system, we expand them in spherical harmonics series:
$\Phi_{\textrm{in}}=\sum_{\lambda,\mu}\beta_{\lambda\mu}\ r^{\lambda}\ Y_{\lambda\mu},$ (27)
$\Phi_{\textrm{out}}=\sum_{\lambda,\mu}\gamma_{\lambda\mu}\ r^{-\lambda-1}\ Y_{\lambda\mu},$ (28)
where $\beta_{\lambda\mu}$ and $\gamma_{\lambda\mu}$ are the expansion coefficients. This form of the expansion assures that the velocity field vanishes outside the impurity, i.e, $\Phi_{\textrm{out}}(r\rightarrow\infty)=0$. At the edge of impurity, the following boundary condition must be satisfied:
$\left.\Phi_{\rm in}\right|_{r=R_{0}}$ $=\left.\Phi_{\rm out}\right|_{r=R_{0}},$ (29)
$\rho_{\rm in}\left(\left.\frac{\partial\Phi_{\rm in}}{\partial r}\right|_{r=R_{0}}-\frac{dR}{dt}\right)$ $=\rho_{\rm out}\left(\left.\frac{\partial\Phi_{\rm out}}{\partial r}\right|_{r=R_{0}}-\frac{dR}{dt}\right).$ (30)
For the shape parametrization (2) we have
$\frac{dR}{dt}=R_{0}\sum_{\lambda,\mu}\dot{\alpha}_{\lambda\mu}\ Y_{\lambda\mu}(\theta,\phi),$ (31)
and the boundary condition reveals expressions for $\beta_{\lambda\mu}$ and $\gamma_{\lambda\mu}$ as:
$\beta_{\lambda\mu}$ $=\gamma_{\lambda\mu}R_{0}^{-2\lambda-1},$ (32)
$\gamma_{\lambda\mu}$ $=R_{0}^{\lambda+3}\frac{\kappa-1}{\kappa\lambda+(\lambda+1)}\dot{\alpha}_{\lambda\mu}$ (33)
with $\kappa=\rho_{\textrm{in}}/\rho_{\textrm{out}}$. Using the relation (divergence theorem)
[…]
we can write the kinetic energy of the impurity as
[…]
Next, substituting the expansions (27)-(28), we obtain the formula :
$T=\frac{1}{2}m_{n}\rho_{\textrm{out}}(\kappa-1)^{2}R_{0}^{5}\sum_{\lambda,\mu}\frac{1}{\kappa\lambda+(\lambda+1)}|\dot{\alpha}_{\lambda\mu}|^{2}.$ (36)
Comparing it with the general formula for the kinetic energy
$T=\frac{1}{2}\sum_{\lambda,\mu}M_{\lambda}|\dot{\alpha}_{\lambda\mu}|^{2}$ (37)
we can identify the associated mass parameters
$M_{\lambda}=m_{n}\rho_{\textrm{out}}(\kappa-1)^{2}R_{0}^{5}\frac{1}{\kappa\lambda+(\lambda+1)}.$ (38)
For the quadruple mode ($\lambda=2$) it reduces to
$M_{2}=m_{n}\rho_{\textrm{out}}(\kappa-1)^{2}\frac{R_{0}^{5}}{2\kappa+3}.$ (39)
The radius $R_{0}$ can be related to the effective mass $M_{\textrm{CM}}$ of the impurity
$R_{0}=\Bigg{(}\frac{3M_{\textrm{CM}}}{4\pi m_{n}\rho_{\textrm{in}}}\Bigg{)}^{1/3}.$ (40)
Finally, we relate the quadrupole moment to the deformation parameter
$Q_{20}=C\alpha_{20},\quad C=\sqrt{\frac{5}{16\pi}}\frac{M_{\textrm{CM}}}{m_{n}}R_{0}^{2}.$ (41)
Then, mass parameter for quadrupole moment oscillations $M_{Q_{20}}$ reads:
$M_{Q_{20}}=\frac{M_{2}}{C^{2}},$ (42)
which gives Eq. (26) from the main text.
Appendix B
Here, we provide a tabularized version of the data presented in Figs 3, 4c, and 6.
$\bar{\rho}$ $\rho_{Bn}$ $M_{CM}$ (A=1fm) $M_{CM}$ (A=10 fm) $\gamma_{\textrm{CM}}$ (A=10 fm) $M_{Q_{20}}$ $\hbar\omega_{Q_{20}}$
[…]
0.0023 0.0016 147.17 152.09 5.05e-06 0.104 21.7
0.0058 0.0045 134.63 142.78 2.02e-05 0.107 22.2
0.0104 0.0084 161.81 171.63 6.03e-05 0.0878 17.2
0.0148 0.0120 151.49 162.52 8.74e-05 0.0870 17.9
0.0187 0.0152 170.91 186.30 7.77e-05 0.0777 16.2
0.0237 0.0193 163.46 180.44 1.38e-04 0.0756 16.1
0.0267 0.0217 162.98 184.11 1.70e-04 0.0719 15.3
0.0300 0.0244 165.54 185.64 2.05e-04 0.0675 14.5
0.0338 0.0276 160.94 178.12 2.74e-04 0.0655 15.6
0.0428 0.0351 147.19 160.73 3.37e-04 0.0567 12.3
0.0510 0.0422 147.15 163.48 5.24e-04 0.0508 11.3
Table 1: The consecutive columns describe the layer of the inner crust of density $\bar{\rho}$ with bulk neutron density $\rho_{Bn}$. We provide the effective mass $M_{\textrm{eff}}$ of a zirconium nucleus for amplitudes of oscillations equal to 1 fm and 10 fm, as well as its susceptibility for deformation $M_{Q_{20}}$. The damping coefficient for A=10fm oscillations is given by $\gamma_{\textrm{CM}}$. The quadrupole mode energy $\hbar\omega_{Q_{20}}$ is depicted in that last column.